OpenStudy (anonymous):

Find the next term of the following sequence. 9, 6, 4, ... I got 3 am I right

OpenStudy (zehanz):

9-6=3 6-4=2 4-3=1 Looks good to me! On the other hand... 9 * 2/3 = 6 6 * 2/3 = 4 4 * 2/3 = 2 2/3

OpenStudy (zehanz):

If it has to be a geometric sequence, then it is 2 2/3